# Improved Trade-offs Between Amortization and Download Bandwidth for Linear HSS

### v1

@@ 1 Introduction
A Homomorphic Secret Sharing (HSS) scheme is a secret sharing scheme that supports computation on top of the shares [Ben86, BGI16b, ${}^{+}$18]. Homomorphic Secret Sharing has been a useful primitive in cryptography, with applications ranging from private information retrieval to secure multiparty computation (see, e.g., [${}^{+}$17, ${}^{+}$18]).
In this work, we focus on information-theoretically secure HSS schemes for the class of degree $d$, $m$-variate polynomials. Suppose $m$ secrets, $x_{1},\ldots,x_{m}$, are shared independently with a $t$-private secret sharing scheme $\mathsf{Share}$111 A $t$-private secret sharing scheme shares a secret $x$ among $s$ servers by computing $s$ shares, $\mathsf{Share}(x)=(y_{1},\ldots,y_{s})$. The $t$-privacy guarantee means that no $t$ of the servers should be able to learn anything about $x$ given their shares. and that server $j$ receives the $m$ shares $y_{k,j}$ for $k\in[m]$. Given a polynomial $f\in{\mathbb{F}}[X_{1},\ldots,X_{m}]$ of degree at most $d$, each server $j$ does some local computation on its shares $y_{k,j}$ for $k\in[m]$ to obtain an output share $z_{j}=\mathsf{Eval}(f,j,(y_{1,j},\ldots,y_{m,j}))$. An output client receives the output shares $z_{1},\ldots,z_{s}$ and runs a recovery algorithm $\mathsf{Rec}$ to obtain $f(x_{1},\ldots,x_{m})=\mathsf{Rec}(z_{1},\ldots,z_{s})$. The HSS scheme $\pi$ is given by the tuple of functions $(\mathsf{Share},\mathsf{Eval},\mathsf{Rec})$; see Definition 2 for a formal definition.
@@
A key parameter of interest in an HSS scheme is the download rate (Definition 4), which is the ratio of the number of bits in $f(x)$ to the number of bits in all of the output shares $z_{j}$. Ideally this rate would be as close to $1$ as possible, because that would mean that the output client does not have to download too much more information than it wishes to compute.
Another parameter of interest is the amortization that the scheme uses. As in previous work [FIKW22, BW23], we consider HSS schemes for low-degree polynomials that amortize over $\ell$ instances of the problem. This means that we have $\ell$ batches of $m$ secrets, $x_{k}^{(i)}$ for $i\in[\ell]$ and $k\in[m]$, and $\ell$ polynomials $f_{1},\ldots,f_{\ell}$. Each of these $m\ell$ secrets is shared independently, as before, but the output shares $z_{j}$ are allowed to depend on all $\ell$ batches. Then the output client is responsible for computing $f_{i}(x_{1}^{(i)},\ldots,x_{m}^{(i)})$ for all $i\in[\ell]$.
@@
In more detail, [FIKW22] showed that for $t$-private, $s$-server linear HSS schemes for $m$-variate degree-$d$ polynomials, the best download rate possible is $1-dt/s$. They achieved this download rate with schemes that had amortization $\ell=\Omega(s\log(s))$. In follow-up work, [BW23] showed that in fact amortization $\ell=\Omega(s\log(s))$ was necessary to acheive the optimal download rate of $1-dt/s$. Their result followed from a characterization optimal-rate linear HSS schemes in terms of a coding theoretic notion they introduce, termed optimal labelweight codes.
In our work, informally, we show that by backing off from the optimal rate by a small amount, we are able to get asymptotic improvements in the amortization required; in some cases we need only $\ell=O(s)$. We obtain this by generalizing the characterization from [BW23] to all HSS schemes, not just optimal ones. We discuss our main results below in more detail.
@@ 1.1 Main Results
    Improved amortization without much loss in rate. The work [BW23] showed that to achieve optimal rate, it was necessary to have amortization $\ell=\Omega(s\log s)$. Leveraging our characterization from Item 1, we give efficient constructions of linear HSS schemes that achieve near-optimal download rate while requiring amortization parameter only $\ell=O(s)$. We compute the parameters of our constructions in practical parameter regimes and show that our schemes achieve an near order-of-magnitude savings in amortization parameter, even for reasonable values of $s,d,m$.
@@
Using the complete characterization of all linear HSS schemes, we construct linear HSS schemes that achieve near-optimal download rate at amortization parameters that are strictly linear in in the number of servers $s$. Through the lens of Theorem 1, this is equivalent to constructing high-labelweight, high-rate linear codes. While the construction of [BW23] use Reed-Solomon codes as a starting point to constructing optimal rate linear HSS schemes, we use two well-studied families of algebraic geometry (AG) codes—Hermitian codes and Goppa codes—and show that such codes yield labelweight codes with attractive parameters for realistic server counts. We loosely summarize these results in Table 1; see Theorems 16, 19 for additional details and formal statements.
@@ 1.2 Technical Overview
Suppose towards a contradiction that for some non-zero $\mathbf{m}\in{\mathbb{F}}^{\ell}$, $\Delta_{\mathcal{L}}(\mathcal{C})\leq t$. Then $\mathbf{m}G\mathbf{x}=\mathbf{c}\mathbf{x}$ (for some non-zero $\mathbf{c}\in{\mathbb{F}}^{n}$) is a linear combination of secrets recoverable by a set of $t$ servers, contradicting the $t$-privacy they were originally secret shared with. ∎
@@ 1.3 Related Work
Though linear HSS schemes are implicit in classical protocols for secure multi-party computation and private information retrieval [Ben86, BGW88, CCD88, CDM00, BF90, BFKR90, CGKS98], the systematic study of HSS was introduced by [${}^{+}$18]. Most HSS schemes reply on cryptographic hardness assumptions [BGI15, DHRW16, BGI16a, BGI16b, FGJI17, ${}^{+}$18, BKS19, ${}^{+}$19, CM21, OSY21, RS21, DIJL23].
In contrast, the HSS schemes presented in this work are information-theoretically secure. The information-theoretic setting was explored in [${}^{+}$18] and was further studied in [FIKW22] and [BW23]; these latter two works are the closest to our work, and we discuss them more below.
The work of [FIKW22] focused the download rate of information-theoretically secure HSS schemes (both linear and non-linear) and proved a tight impossibility result regarding the highest download rate achievable by linear HSS schemes. They paired this with an explicit construction that showed a large amortization parameter is a sufficient condition for a linear HSS scheme to achieve optimal download rate.
@@
Let $s,d,t\in\mathbb{Z}^{+}$ such that $s>dt$. Let $\pi$ be a $t$-private, $s$-server linear HSS scheme for $\mathrm{POLY}_{d,m}({\mathbb{F}})$. Then $\mathsf{DownloadRate}(\pi)\geq 1-dt/s$.
@@
The work [BW23] focuses linear schemes with optimal download rate, meeting the bound showed in [FIKW22]. They proved that in fact a large amortization parameter is necessary for a linear HSS scheme to achieve optimal download rate.
@@
[…]
@@
[…]
@@ 2.1 Homomorphic Secret Sharing
Given a collection of $s$ servers and a function class =$\mathcal{F}=\{f:\mathcal{X}^{m}\to\mathcal{O}\}$, consider a tuple $\pi=(\mathsf{Share},\mathsf{Eval},\mathsf{Rec})$, where $\mathsf{Share}:\mathcal{X}\times\mathcal{R}\to\mathcal{Y}^{s}$, $\mathsf{Eval}:\mathcal{F}\times[s]\times\mathcal{Y}\to\mathcal{Z}^{\ast}$, and $\mathsf{Rec}:\mathcal{Z}^{\ast}\to\mathcal{O}$ as follows [^3] By $\mathcal{Z}^{\ast}$, we mean a vector of some number of symbols from $\mathcal{Z}$.:
@@
Then there exists a matrix $G_{\pi}\in\mathbb{F}^{\ell\times n}$ so that, for all $f\in\mathcal{F}$ and for all secrets $\mathbf{x}\in({\mathbb{F}}^{m})^{\ell}$,
$\mathsf{Rec}(\mathbf{z})=G_{\pi}\mathbf{z}=f(\mathbf{x})=\left[f_{1}(\mathbf{x}^{(1)}),\;f_{2}(\mathbf{x}^{(2)}),\ldots,\;f_{\ell}(\mathbf{x}^{(\ell)})\right]^{\mathsf{T}}$
@@
Let $s,t$ be integers and let $\mathcal{F}$ be a class of functions with input space $\mathcal{X}^{m}$ and output space $\mathcal{O}$. Let $\pi$ be a $s$-server $t$-private HSS for $\mathcal{F}$. Let $z_{i}\in\mathcal{Z}^{n_{i}}$ for $i\in[s]$ denote the output shares.
@@ 3 Equivalence of Linear HSS and Labelweight Codes
[…]
@@
[…]
@@
Let $\ell,t,s,d,m,n$ be integers, with $m\geq d$. Suppose that there exists a linear code $\mathcal{C}\subseteq{\mathbb{F}}^{n}$ with dimension $\ell$ and rate $\ell/n$. Fix a labeling $\mathcal{L}:[n]\to[s]$ so that $\Delta_{\mathcal{L}}(\mathcal{C})\geq dt+1$. Then there exists a $t$-private, $s$-server linear HSS $\pi=(\mathsf{Share},\mathsf{Eval},\mathsf{Rec})$ for $\mathrm{POLY}_{d,m}({\mathbb{F}})^{\ell}$ with download rate $\ell/n$ and amortization parameter $\ell$.
To prove this direction without the strong symmetries of optimal rate linear HSS, we begin with the following lemma.
@@
Let $G_{\pi}$ be a reconstruction matrix. At a high level, Lemma 13 says that any sufficiently large submatrix of $G_{\pi}$, obtained by only considering columns labeled with a sufficiently large subset $\Lambda\subseteq[s]$, must be full-rank. [BW23] proved that such a property held for $G_{\pi}$ in the optimal-rate regime that relied heavily on the fact that, in optimal-rate linear HSS, the output client downloads and equal number of output symbols from each server. The proof of Lemma 13 shows that, perhaps surprisingly, sufficiently large submatrices of $G_{\pi}$ still achieve full-rank even when the output client is allowed to download arbitrary numbers of output symbols from each server.
@@ 4 Linear HSS from Hermitian Codes
In the following sections we take a different approach and base our labelweight codes on well-studied codes from algebraic geometry. In particular, this section constructs a family of linear HSS schemes from Hermitian codes; notably, such schemes achieve asymptotically optimal download rate while requiring an amortization parameter that is only linear in $s$.
@@
Let $\ell,t,s,d,m$ be positive integers and $q$ a prime power satisfying $m\geq d$, $s-dt>0$, and $s=q^{3}$. Then there exists a $t$-private, $s$-server HSS $\pi=\left(\mathsf{Share},\mathsf{Eval},\mathsf{Rec}\right)$ for any non-trivial $\mathcal{F}\subseteq\mathrm{POLY}_{d,m}(\mathbb{F}_{q^{2}})$ with

### v2

@@ 1 Introduction
A Homomorphic Secret Sharing (HSS) scheme is a secret sharing scheme that supports computation on top of the shares [Ben86, BGI16b, BGI+18]. Homomorphic Secret Sharing has been a useful primitive in cryptography, with applications ranging from private information retrieval to secure multiparty computation (see, e.g., [BCG+17, BGI+18]).
In this work, we focus on information-theoretically secure HSS schemes for the class of degree $d$, $m$-variate polynomials. Suppose $m$ secrets, $x_{1},\ldots,x_{m}$, are shared independently with a $t$-private secret sharing scheme $\mathsf{Share}$111 A $t$-private secret sharing scheme shares a secret $x$ among $s$ servers by computing $s$ shares, $\mathsf{Share}(x)=(y_{1},\ldots,y_{s})$. The $t$-privacy guarantee means that no $t$ of the servers should be able to learn anything about $x$ given their shares. and that server $j$ receives the $m$ shares $y_{k,j}$ for $k\in[m]$, $j\in[s]$. Denote by $\mathrm{POLY}_{d,m}({\mathbb{F}})\subseteq\mathbb{F}[X_{1},\ldots,X_{m}]$ an arbitrary set of degree $d$, $m$-variate polynomials. Given a polynomial $f\in\mathrm{POLY}_{d,m}({\mathbb{F}})$, each server $j$ does some local computation on its shares $y_{k,j}$ for $k\in[m]$ to obtain an output share $z_{j}=\mathsf{Eval}(f,j,(y_{1,j},\ldots,y_{m,j}))$. An output client receives the output shares $z_{1},\ldots,z_{s}$ and runs a recovery algorithm $\mathsf{Rec}$ to obtain $f(x_{1},\ldots,x_{m})=\mathsf{Rec}(z_{1},\ldots,z_{s})$. The HSS scheme $\pi$ is given by the tuple of functions $(\mathsf{Share},\mathsf{Eval},\mathsf{Rec})$; see Definition 2 for a formal definition.
@@
A key parameter of interest in an HSS scheme is the download rate (Definition 4), which is the ratio of the number of bits in $f(x)$ to the number of bits in all of the output shares $z_{j}$. Ideally this rate would be as close to $1$ as possible; i.e., the output client does not download much more information than it wishes to compute.
Another parameter of interest is amortization. As in previous work [FIKW22, BW23], we consider HSS schemes for low-degree polynomials that amortize over $\ell$ instances of the problem. This means that we have $\ell$ batches of $m$ secrets, $x_{k}^{(i)}$ for $i\in[\ell]$ and $k\in[m]$, and $\ell$ polynomials $f_{1},\ldots,f_{\ell}$. Each of these $m\ell$ secrets is shared independently as before, but the output shares $z_{j}$ are allowed to depend on all $\ell$ batches. Then the output client is responsible for computing $f_{i}(x_{1}^{(i)},\ldots,x_{m}^{(i)})$ for all $i\in[\ell]$.
@@
In more detail, [FIKW22] showed that for $t$-private, $s$-server linear HSS schemes for $m$-variate degree-$d$ polynomials, the best download rate possible is $1-dt/s$. They achieved this download rate with schemes that had amortization $\ell=\Omega(s\log(s))$. In follow-up work, [BW23] showed that in fact amortization $\ell=\Omega(s\log(s))$ was necessary to achieve the optimal download rate of $1-dt/s$. Their result followed from a characterization optimal-rate linear HSS schemes in terms of a coding theoretic notion they introduce, termed optimal labelweight codes.
In our work, informally, we show that by backing off from the optimal rate by a small amount, we are able to get asymptotic improvements in the amortization required; in some cases we need only $\ell=O(s)$. We obtain this by generalizing the characterization from [BW23] to all HSS schemes, not just optimal rate ones. We discuss our main results below in more detail.
@@ 1.1 Main Results
    Improved amortization without much loss in rate. The work [BW23] showed that to achieve optimal rate, it is necessary to have amortization $\ell=\Omega(s\log s)$. Leveraging our characterization from Item 1, we give efficient constructions of linear HSS schemes that achieve near-optimal download rate while requiring amortization parameter only $\ell=O(s)$. We compute the parameters of our constructions in practical parameter regimes and show that our schemes achieve a near order-of-magnitude savings in amortization parameter, even for reasonable values of $s,d,m$.
@@
Using the complete characterization of all linear HSS schemes, we construct linear HSS schemes that achieve near-optimal download rate at amortization parameters that are strictly linear in in the number of servers $s$. While the construction of [BW23] uses Reed-Solomon codes as a starting point to construct optimal rate linear HSS schemes, we use two well-studied families of algebraic geometry (AG) codes—Hermitian codes and Goppa codes.
Through the lens of Theorem 1, finding low amortization, high download rate linear HSS schemes is equivalent to constructing low dimension, high rate labelweight codes; that is, we wish for code dimension $\ell$ to be minimized and for code rate $\ell/n$ to be maximized. Observe that for any code dimension $\ell$, code rate is maximized when $n$ is minimized. Since $n\geq s$, we see that this occurs when setting $n=s$. In this setting, it immediately follows that the surjective labeling function $\mathcal{L}:[n]\to[s]$ can be assumed to be the identity map without loss of generality.
Therefore, our constructions use the trivial labeling scheme where $n=s$ and $\mathcal{L}:[n]\to[s],x\mapsto x$ is the identity function. Such labelweight codes, though straightforward, yield linear HSS schemes with attractive parameters for realistic server counts. Furthermore, their intuitive construction underscores the fundamental relationship between classical codes and linear HSS. We loosely summarize these results in Table 1; see Theorems 16, 19 for additional details and formal statements.
@@ 1.2 Technical Overview
Suppose towards a contradiction that for some non-zero $\mathbf{m}\in{\mathbb{F}}^{\ell}$, $\Delta_{\mathcal{L}}(\mathbf{m}G)\leq t$. Then $\mathbf{m}G\mathbf{x}=\mathbf{c}\mathbf{x}$ (for some non-zero $\mathbf{c}\in{\mathbb{F}}^{n}$) is a linear combination of secrets recoverable by a set of $t$ servers, contradicting the $t$-privacy they were originally secret shared with. ∎
@@ 1.3 Related Work
Though linear HSS schemes are implicit in classical protocols for secure multi-party computation and private information retrieval [Ben86, BGW88, CCD88, CDM00, BF90, BFKR90, CGKS98], the systematic study of HSS was introduced by [BGI+18]. Most HSS schemes reply on cryptographic hardness assumptions [BGI15, DHRW16, BGI16a, BGI16b, FGJI17, BGI+18, BKS19, BCG+19, CM21, OSY21, RS21, DIJL23].
In contrast, the HSS schemes presented in this work are information-theoretically secure. The information-theoretic setting was explored in [BGI+18] and was further studied in [FIKW22] and [BW23]; these latter two works are the closest to our work, and we discuss them more below.
The work of [FIKW22] focused on the download rate of information-theoretically secure HSS schemes (both linear and non-linear) and proved a tight impossibility result regarding the highest download rate achievable by linear HSS schemes. They paired this with an explicit construction that showed a large amortization parameter is a sufficient condition for a linear HSS scheme to achieve optimal download rate.
@@
Let $s,d,t\in\mathbb{Z}^{+}$ such that $s>dt$. Let $\pi$ be a $t$-private, $s$-server linear HSS scheme for $\mathrm{POLY}_{d,m}({\mathbb{F}})$. Then $\mathsf{DownloadRate}(\pi)\leq 1-dt/s$.
@@
The work [BW23] focused on linear schemes with optimal download rate, meeting the bound showed in [FIKW22]. They proved that in fact a large amortization parameter is necessary for a linear HSS scheme to achieve optimal download rate.
@@
There exists a $t$-private, $s$-server linear HSS scheme for $\text{POLY}_{d,m}(\mathbb{F})$ with download rate $(s-dt)/s$ and amortization parameter $\ell$ only if $\ell=j(s-dt)$ for some $j\in\mathbb{Z}+$ satisfying
@@
There exists a $t$-private, $s$-server linear HSS scheme for $\text{POLY}_{d,m}(\mathbb{F})$ with download rate $(s-dt)/s$ and amortization parameter $\ell$ if and only if there exists a linear code $\mathcal{C}\subseteq\mathbb{F}^{n}$ with information rate $(s-dt)/s$ and dimension $\ell$; and surjection $L:[n]\to[s]$ such that $\Delta_{L}(\mathcal{C})\geq dt+1$.
@@ 2.1 Homomorphic Secret Sharing
Given a collection of $s$ servers and a function class =$\mathcal{F}=\{f:\mathcal{X}^{m}\to\mathcal{O}\}$, consider a tuple $\pi=(\mathsf{Share},\mathsf{Eval},\mathsf{Rec})$, where $\mathsf{Share}:\mathcal{X}\times\mathcal{R}\to\mathcal{Y}^{s}$, $\mathsf{Eval}:\mathcal{F}\times[s]\times\mathcal{Y}\to\mathcal{Z}^{\ast}$, and $\mathsf{Rec}:\mathcal{Z}^{\ast}\to\mathcal{O}$ are as follows [^3] By $\mathcal{Z}^{\ast}$, we mean a vector of some number of symbols from $\mathcal{Z}$.:
@@
Then there exists a matrix $G_{\pi}\in\mathbb{F}^{\ell\times n}$ so that, for all $\mathbf{f}\in\mathcal{F}$ and for all secrets $\mathbf{x}\in({\mathbb{F}}^{m})^{\ell}$, there exists some $\mathbf{z}\in\mathbb{F}^{n}$ such that
$\mathsf{Rec}(\mathbf{z})=G_{\pi}\mathbf{z}=\mathbf{f}(\mathbf{x})=\left[\mathbf{f}_{1}(\mathbf{x}^{(1)}),\;\mathbf{f}_{2}(\mathbf{x}^{(2)}),\ldots,\;\mathbf{f}_{\ell}(\mathbf{x}^{(\ell)})\right]^{\mathsf{T}}.$
@@
Let $s,t$ be integers and let $\mathcal{F}$ be a class of functions with input space $\mathcal{X}^{m}$ and output space $\mathcal{O}$. Let $\pi$ be an $s$-server $t$-private HSS for $\mathcal{F}$. Let $z_{i}\in\mathcal{Z}^{n_{i}}$ for $i\in[s]$ denote the output shares.
@@ 3 Equivalence of Linear HSS and Labelweight Codes
Let $\ell,t,s,d,m,n$ be integers, with $m\geq d$, $\ell\leq n$. There exists a $t$-private, $s$-server $\mathbb{F}$-linear HSS $\pi=\left(\mathsf{Share},\mathsf{Eval},\mathsf{Rec}\right)$ for any non-trivial $\mathcal{F}\subseteq\text{POLY}_{d,m}(\mathbb{F})^{\ell}$, with download rate $\textsf{DownloadRate}(\pi)=\ell/n$, if and only if there exists a linear code $\mathcal{C}\subseteq{\mathbb{F}}^{n}$ with rate $\textsf{DownloadRate}(\pi)$ and a labeling $\mathcal{L}:[n]\to[s]$ so that $\Delta_{\mathcal{L}}(\mathcal{C})\geq dt+1$.
@@
Let $\ell,t,s,d,m,n$ be integers, with $m\geq d$, $\ell\leq n$. Suppose there exists a $t$-private, $s$-server $\mathbb{F}$-linear HSS $\pi=\left(\mathsf{Share},\mathsf{Eval},\mathsf{Rec}\right)$ for any non-trivial (see Definition 6) $\mathcal{F}\subseteq\text{POLY}_{d,m}(\mathbb{F})^{\ell}$, with download rate $\textsf{DownloadRate}(\pi)=\ell/n$. Then there exists a linear code $\mathcal{C}\subseteq{\mathbb{F}}^{n}$ with rate $\textsf{DownloadRate}(\pi)$ and a labeling $\mathcal{L}:[n]\to[s]$ so that $\Delta_{\mathcal{L}}(\mathcal{C})\geq dt+1$.
@@
Let $\ell,t,s,d,m,n$ be integers, with $m\geq d$. Suppose that there exists a linear code $\mathcal{C}\subseteq{\mathbb{F}}^{n}$ with dimension $\ell$ and rate $\ell/n$. Suppose there exists a labeling $\mathcal{L}:[n]\to[s]$ so that $\Delta_{\mathcal{L}}(\mathcal{C})\geq dt+1$. Then there exists a $t$-private, $s$-server linear HSS $\pi=(\mathsf{Share},\mathsf{Eval},\mathsf{Rec})$ for $\mathrm{POLY}_{d,m}({\mathbb{F}})^{\ell}$ with download rate $\ell/n$ and amortization parameter $\ell$.
The main ingredient in proving this direction without the strong symmetries of optimal rate linear HSS is the following lemma, which neatly generalizes the results of Lemma 13 and Corollaries 14, 15 of [BW23].
@@
Let $G_{\pi}$ be a reconstruction matrix. At a high level, Lemma 13 says that any sufficiently large submatrix of $G_{\pi}$, obtained by only considering columns labeled with a sufficiently large subset $\Lambda\subseteq[s]$, must be full-rank. [BW23] proved that such a property held for $G_{\pi}$ in the optimal-rate regime that relied heavily on the fact that, in optimal-rate linear HSS, the output client downloads an equal number of output symbols from each server. This is equivalent to requiring that the sets $\mathcal{L}^{-1}(y):=\{x\in[n]:\mathcal{L}(x)=y\}$ be the same size for all $y\in[s]$. The proof of Lemma 13 shows that, perhaps surprisingly, sufficiently large submatrices of $G_{\pi}$ still achieve full-rank even when the output client is allowed to download arbitrary numbers of output symbols from each server.
@@ 4 Linear HSS from Hermitian Codes
In the following sections we take a different approach. We derive our labelweight codes straightforwardly from well-studied algebraic geometric codes: we set the number of servers $s$ equal to the block length $n$ of the codes and label each coordinate by the identity function $\mathcal{L}:[n]\to[s],x\mapsto x$. In this setting, labelweight is equivalent to Hamming weight. We justify this straightforward approach in subsection 2 of Section 1.1.
This section constructs a family of linear HSS schemes from Hermitian codes; notably, such schemes achieve asymptotically optimal download rate while requiring an amortization parameter that is only linear in $s$.
@@
Let $\ell,t,s,d,m$ be positive integers and $q$ a prime power satisfying $m\geq d$, $s-dt>0$, and $s=q^{3}$. Then there exists an explicit $t$-private, $s$-server HSS $\pi=\left(\mathsf{Share},\mathsf{Eval},\mathsf{Rec}\right)$ for any non-trivial $\mathcal{F}\subseteq\mathrm{POLY}_{d,m}(\mathbb{F}_{q^{2}})$ with